OpenStudy (anonymous):

y-x=15 5x-2y=525 help medal

OpenStudy (jackthegreatest):

y= x + 15

OpenStudy (jackthegreatest):

plug this in for y in the equation 5x-2y=525

OpenStudy (jackthegreatest):

and u get 5x- 2 (x + 15)= 525

OpenStudy (anonymous):

thats two different equations though

OpenStudy (jackthegreatest):

5x -2x - 30 = 525

OpenStudy (anonymous):

trying to get x and y

OpenStudy (jackthegreatest):

i k

OpenStudy (jackthegreatest):

first find y

OpenStudy (jackthegreatest):

then plug it in to get x

OpenStudy (jackthegreatest):

3x -30= 525

OpenStudy (jackthegreatest):

3x = 565

OpenStudy (jackthegreatest):

* 555

OpenStudy (jackthegreatest):

x = 185

OpenStudy (jackthegreatest):

plug that back into y = x + 15

OpenStudy (jackthegreatest):

wat is y then @imdestroyer69

OpenStudy (anonymous):

200

OpenStudy (jackthegreatest):

correcto : )

OpenStudy (anonymous):

thk

OpenStudy (jackthegreatest):

welcome